# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

### v1

@@ 1.2 Contributions
    We derive a novel algorithm Action Gradient MCTS (AGMCTS) that combines the MIS tree with action gradients. This algorithm is applicable to both MDPs and POMDPs. We show several algorithmic improvements and heuristics to make it more efficient.
@@
    We compare AGMCTS to PFT-DPW and POMCPOW in a simulated POMDP scenario at several dimensions and different noise levels in the domain knowledge. AGMCTS had the best overall performance on a per-simulation budget.
2 Background
MDPs.
An MDP is the tuple $\langle\mathcal{S},\mathcal{A},p_{T},r,\gamma,L,b_{0}\rangle$. $\mathcal{S},\mathcal{A}$ are the state and action spaces. The transition model $p_{T}(s^{\prime}\mid s,a)$ models the probability of arriving at state $s^{\prime}$ when taking action $a\in\mathcal{A}$ at state $s\in\mathcal{S}$. The reward function $r(s,a,s^{\prime})\in\mathbb{R}$ gives the immediate reward of transitioning from state $s$ to $s^{\prime}$ by taking action $a$. We define the reward of $(s,a)$ as the expected reward over outcome states $r(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$. $\gamma\in(0,1]$ is the discount factor. The MDP starts at time 00 and terminates after $L\in\mathbb{N}\cup\{\infty\}$ steps. The initial state is drawn from the distribution $s_{0}\sim b_{0}$.
A timed policy at time $i$, denoted as $\pi_{i}$, is a mapping $\pi_{i}\colon\mathcal{S}\to\mathcal{A}$. A policy $\pi$ is a collection of timed policies from the start time until the horizon: $\pi\triangleq(\pi_{t})_{t=1}^{t=L}$. The value function of a policy $\pi$ at time $t$ is the expected sum of discounted rewards until the horizon: $V_{t}^{\pi}(s_{t})\triangleq\mathbb{E}_{s_{t+1:L}\mid s_{t},\pi}[\sum_{i=t}^{L}\gamma^{i-t}r(s_{i},\pi_{i}(s_{i}),s_{i+1})]$. The action-value function is often used as an intermediate step in the estimation of the value function, and is defined as $Q_{t}^{\pi}(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime})]$, where state $s^{\prime}$ is at the proceeding time of $s$. The goal of planning is to compute a policy that maximizes the value function.
POMDPs.
A POMDP adds $\langle\mathcal{O},p_{O}\rangle$ to the MDP tuple. The observation space is $\mathcal{O}$, and the observation model $p_{O}(o\mid s)$ is the conditional probability of receiving an observation $o\in\mathcal{O}$ at $s\in\mathcal{S}$. A history at time $t$, $H_{t}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1},o_{t})$, is defined as a sequence of the starting belief, followed by actions taken and observations received until $t$. Because of the partial observability, the agent has to maintain a probability distribution of the current state given past actions and observations, known as the belief. The belief space, denoted in $\mathcal{B}$, is the space of all probability distributions over $\mathcal{S}$. The belief at time $t$ is defined as the conditional distribution of the state $b_{t}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t})$, given the history. We denote $H_{t}^{-}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1})$ for the same history without the last measurement, and the propagated belief $b_{t}^{-}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t}^{-})$. It can be shown that optimal decision-making can be made given the belief, instead of considering the entire history [?]. Since the belief fulfills the Markov property, a POMDP is an MDP on the belief space, commonly referred to as belief MDP. Policies, the value function and the action-value function, extend from MDPs to POMDPs by equivalent definitions on $b\in\mathcal{B}$ as states. As an example, $r(b,a)\triangleq\mathbb{E}_{b^{\prime}\mid b,a}[r(b,a,b^{\prime})]$. In the general case, the belief-reward function may be any function of beliefs, f.e. entropy, and this is called $\rho$-POMDP [?]. Often, the belief reward has an underlying state-reward, in the form $r(b,a,b^{\prime})=\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]$. We refer to this case as a state-reward POMDP.
Importance Sampling (IS).
The Monte-Carlo (MC) estimator of an expectation, denoted with $\hat{\mathbb{E}}$ or f.e. $\hat{V}_{t}^{\pi}$, is its approximation with a finite number of samples. Importance Sampling (IS) is a technique in MC estimation where a proposal distribution $q$ is used to sample from instead of $p$. The resulting estimator for $f(x)$ is $\hat{f}=\frac{1}{N}\sum_{i=1}^{N}\frac{p(x^{i})}{q(x^{i})}f(x^{i})$. This estimator is consistent if $q(x)=0\Rightarrow p(x)=0$. With a proper choice of $q$ the variance of the estimator can be reduced, or we can calculate estimates for a target $p$ which is difficult to sample directly. Commonly the sampled weights $w_{i}=p(x^{i})/q(x^{i})$ are normalized to sum to $1$, giving the self-normalized IS estimator [?, 1.3.2].
Particle Beliefs.
Often in continuous settings we cannot maintain an exact belief. Hence, particle filters are used instead to represent a belief non parametrically [?]. We denote particle beliefs with $\bar{b}$. The particle belief of $J$ particles is represented by a set of state-weight pairs $\bar{b}=\{(s^{j},w^{j})\}_{j=1}^{J}$, and is defined as the discrete distribution $\mathbb{P}(s\mid\bar{b})={\sum_{j=1}^{J}{w^{j}\cdotp\delta(s-s^{j})}}/{\sum_{j=1}^{J}{w^{j}}}$. The basic particle filtering algorithm, Sequential Importance Sampling (SIS) [?, 1.3.2], recursively updates and normalizes the importance weights of the particles based on the observation likelihood. Sequential Importance Resampling (SIR) adds resampling of state particles after the weight updates, to avoid weight degeneracy [?]. The particle belief MDP (PB-MDP) was considered in [?], where the belief space is explicitly the space of ordered particle beliefs $\bar{b}$ of $J$ particles. An ordered particle belief is the ordered set $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$.
Multiple Importance Sampling (MIS).
MIS [?] is a method to combine the advantages of several sampling methods. It is defined as $F=\sum_{i=1}^{n}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}w_{i}(x^{i,j})\frac{f(x^{i,j})}{q_{i}(x^{i,j})}$, which is a weighted average of $n$ IS estimators. If the weights $w_{i}$ satisfy $\sum_{i=1}^{n}w_{i}(x)=1$ whenever $f(x)\neq 0$, and $w_{i}(x)=0$ whenever $q_{i}(x)=0$, then $F$ is unbiased. The choice of $w_{i}$ may depend on the samples $x^{i,j}$. The weighting strategy $w_{i}(x)=\frac{n_{i}q_{i}(x)}{\sum_{k}n_{k}q_{k}(x)}$, named the balance heuristic, has its variance provably bounded from the optimal weighting strategy.
Monte Carlo Tree Search (MCTS).
MCTS is an algorithm used to quickly explore large state spaces [?]. It iteratively repeats four steps to build a search tree that approximates the action-values, using a best-first strategy: (1) Selection Starting from the root node, descend recursively until a node to which children can be added is found; (2) Expansion A new node is added as a child, according to action expansion strategy; (3) Simulation A simulation is run from the new node according to the rollout (default) policy; (4) Backpropagation The simulation result is backpropagated through the selected nodes to update the action-values. Often UCT is used at selection to balance between exploration of new actions and exploitation of promising ones [?]. Double Progressive Widening (DPW) [?] is a technique to limit the branching factor from being infinite in continuous settings. The number of children of a node is artificially limited to $kN^{\alpha}$ for $N$ visitations, for hyperparameters $k$ and $\alpha$. The progressive widening is double due to being applied both to the branching of actions, and branching of states.
Stochastic Gradient Optimization (SGO).
SGO is a widely used technique for efficient optimization of differentiable functions [?]. Gradient optimizers tremendously rose in popularity to their applications in deep neural networks. Adam [?] is a widely used algorithm, with heuristically adaptive step size based on first and second moment estimates of the gradient. Adam has been shown to perform well with noisy gradient estimates, and requires minimal tuning to different problems.
3 Methodology
3.1 Action-Values with Importance Sampling
Figure 2: Recursive definitions of the MIS tree, for the particle belief MDP (PB-MDP) case, in which states are particle beliefs. State nodes are circles and action nodes are squares. The state $\bar{b}$ has a value estimate $\hat{V}(\bar{b})$ and visitation count $n(\bar{b})$. The action node $(\bar{b},a)$ has an immediate reward estimate $\hat{r}(\bar{b},a)$, a future value estimate $\hat{V}_{f}(\bar{b},a)$, and visitation count $n(\bar{b},a)$. The action-value estimate is defined via $\hat{Q}(\bar{b},a)\triangleq\hat{r}(\bar{b},a)+\gamma\hat{V}_{f}(\bar{b},a)$. During action update we update $(\bar{b},a)$ to $(\bar{b},a^{\prime})$. We highlight that in the PB-MDP, we also store the propagated belief $\bar{b}^{-}$. The child state node has all properties as $\bar{b}$, and stores the proposal action $a_{prop}(\bar{b}^{-})$. During action update (11) we do $\omega^{\prime}={p_{T}(\bar{b}^{-}\mid\bar{b},a^{\prime})}({p_{T}(\bar{b}^{-}\mid\bar{b},a_{prop}(\bar{b}^{-}))})^{-1}$. In this example, $a_{prop}(\bar{b}^{-})=a$.
In an MDP setting, when updating action $a$ to $a^{\prime}$, we would like to estimate the new action-value $Q_{t}^{\pi}(s,a^{\prime})$ by reusing the information that was used for estimating $Q_{t}^{\pi}(s,a)$. Otherwise, updating actions would require us to theoretically recompute the entire subtree, which would render the algorithm very inefficient. Since this process might occur several times for each action branch, in the general case each child state may have been generated from a different proposal action. Hence, we have to account for an MIS scenario in order to estimate $Q_{t}^{\pi}(s,a^{\prime})$.
Assuming that $p_{T}(s^{\prime}\mid s,a^{\prime})$ is a non-strict subset of the support of $p_{T}(s^{\prime}\mid s,a)$, we can formally write
$\textstyle Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))].$ (1)
For POMDPs, the posterior belief transition likelihood is generally intractable. However, by using the law of total expectation, we introduce a dependence of the posterior belief $b^{\prime}$ on the propagated $b^{-}$. This makes only $b^{-}$ directly dependent on $a$, which allows us to introduce an importance ratio over the propagated belief rather than posterior:
$\textstyle Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\textstyle[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))].$ (2)
This is known as the propagated belief trick [?]. We can see that (1) and (2) are analogous, except for the importance ratio being between propagated beliefs in the POMDP case. The propagated belief likelihood is rather simple to calculate for appropriate choice of the belief space.
Lemma 1.
Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR. Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and:
$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (3)
We refer to the appendix for all proofs
3.2 Action Adaptive MIS Trees
We now present our definitions of the novel MIS tree, a recursive tree structure with MIS value estimates, that allows sharing value information between sibling actions. Similarly to belief trees, the tree has levels alternating with states and action nodes, starting from the root state node. Yet in contrast, the MIS tree supports replacing an action node with a new action, and updating the value estimates of the new action and its parent state node based on previous samples. We later introduce action updates via gradient-based optimization.
For the following discussion, we present notations for recursively defining estimators at a state node $s$ in the MIS tree, or some action node $(s,a)$. The generalization to POMDPs is done by taking ordered particle belief states $\bar{b}$, and the following definitions are analogous, illustrated in Figure 2.
Denote the children actions of $s$ as $C(s)=\{a^{i}\}_{i=1}^{\lvert{C(s)}\rvert}$, and the children states of $(s,a)$ as $C(s,a)=\{s^{\prime,i}\}_{i=1}^{M}$, where $M\triangleq\lvert{C(s,a)}\rvert$. The visitation counts of state and action nodes are denoted $n(s)$ and $n(s,a)$ respectively. We denote the proposal action $a_{prop}(s)$, the action that was used to sample $s$, as possibly $a_{prop}(s^{\prime})\neq a$ for $s^{\prime}\in C(s,a)$. We assign each action node $(s,a)$ an action-value estimate $\hat{Q}(s,a)$ that estimates $Q_{t}^{\pi}(s,a)$, assuming $s$ is a state at time $t$. Compared to traditional MCTS algorithms, we also estimate the value function at each state node. We define it as a weighted mean:
$\textstyle\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}).$ (4)
The following two relationships between visitation counts are maintained in MCTS:
$\textstyle n(s)=\sum_{C(s)}n(s,a^{i}),\;\textstyle n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (5)
where $n(s^{\prime,i})_{+1}\triangleq n(s^{\prime,i})+1$. The asymmetry is the result of initializing state nodes to $n(s)=0$ when assigned a rollout value, whereas action nodes are initialized to $n(s,a)=1$. This allows value estimates to be updated completely when they are visited again, and for their parent’s action-value estimate to remember the initial rollout estimate.
We split $\hat{Q}(s,a)$ into two estimators: (1) Immediate reward estimate $\hat{r}(s,a)$, for $\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$; (2) Future value estimate $\hat{V}_{f}(s,a)$ for $\mathbb{E}_{s^{\prime}\mid s,a}[V^{\pi}(s^{\prime})]$ of the proceeding time, such that if $s$ is a state at time $t$, then $\hat{V}_{f}(s,a)$ estimates the expected $V_{t+1}^{\pi}(s^{\prime})$. Together they define $\hat{Q}(s,a)\triangleq\hat{r}(s,a)+\gamma\hat{V}_{f}(s,a)$.
Generally in MIS, each proposal distribution $q_{i}$ has $n_{i}$ samples. In our situation, we assume that each of $C(s,a)$ may have been sampled from a different action, has its own visitation counter and an associated single value estimate $V(s^{\prime,i})$. Therefore, we adapt the most naive unbiased weighting, that is to give equal weight to all samples, and weight the posterior states by the visitation counts. Hence, the naive MIS estimator for the future value is
$\textstyle\hat{V}_{f}(s,a)=(n(s,a))^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (6)
where $\omega^{i}={p_{T}(s^{\prime,i}\mid s,a)}({p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i}))})^{-1}$. The naive MIS estimator for $\hat{r}(s,a)$ is identical, with $\hat{V}(s^{\prime,i})$ replaced by $r(s,a,s^{\prime,i})$.
In practice, this estimator has very bad performance. On the other hand, the balance heuristic requires $O(M)$ time to update [?], which can be quite expensive. We found a simple solution which is to self-normalize the weights. In regular IS, it is known to reduce the variance of the estimator at the expense of introducing bias, and asymptotically converges to the true estimate under weak assumptions [?, 1.3.2]. We define self-normalized MIS (SN-MIS) by
$\textstyle\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (7)
$\textstyle\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (8)
and again the SN-MIS estimator for $\hat{r}(s,a)$ is defined by replacing $\hat{V}(s^{\prime,i})$ with $r(s,a,s^{\prime,i})$.
3.3 Updating MIS Trees in MCTS
We now discuss how to update the SN-MIS estimators in the steps of the MCTS routine. We highlight that for any other MIS weighting heuristic, a similar analysis would follow.
Currently, we do not address selection, and we assume the UCT criterion is not altered. Simulation is unaffected, as it is determined only by the rollout policy. Hence, we analyze the different subcases of backpropagation and expansion steps. We add a new step to MCTS named action update, where we update action node $(s,a)$ to $(s,a^{\prime})$ with the new action $a^{\prime}$, and we show its update equations. In general, action update can be applied at every point of MCTS.
For the recursive analysis, we assume backpropagated values for the children nodes are already updated, and at each step we only update an action node $(s,a)$ and its parent node $s$. The updated values are denoted with a prime, and we assume that we have access to the values before the update. The transition likelihood terms $w_{i}$ are usually the most expensive to calculate, hence it is important to cache them, and recalculate the numerator of $w_{i}$ only when necessary. The derivations of the update equations, alongside numerically stable forms with log-likelihoods are given in the appendix.
@@ 3.3 Updating MIS Trees in MCTS
Let $s^{\prime,i}$ be an updated node. Hence, we updated $n(s^{\prime,i})$ to $n^{\prime}(s^{\prime,i})$ and $\hat{V}(s^{\prime,i})$ to $\hat{V}^{\prime}(s^{\prime,i})$. We update $n(s,a)$ by (5) and perform
$\textstyle\eta^{\prime}=\eta+\omega^{i}(n{{}^{\prime}}(s^{\prime,i})-n(s^{\prime,i})),$ (9)
$\textstyle\hat{V}_{f}^{\prime}(s,a)=(\eta^{\prime})^{-1}(\eta\hat{V}_{f}(s,a)+$
$\textstyle\omega^{i}(n{{}^{\prime}}(s^{\prime,i})\hat{V}^{\prime}(s^{\prime,i})-n(s^{\prime,i})\hat{V}(s^{\prime,i})).$ (10)
@@
This case is the same as action backpropagation, where the updated index is a new summand at index $i=M+1$. The same equations apply with $n(s^{\prime,i})=0$.
@@
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$\textstyle(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))^{-1},$ (11)
and recalculate (7) and (8) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
@@
$\textstyle\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (12)
@@
Let $(s,a^{i})$ be an updated node. Hence, we updated $n(s,a^{i})$ to $n^{\prime}(s,a^{i})$ and $\hat{Q}(s,a^{i})$ to $\hat{Q}^{\prime}(s,a^{i})$. We update $n^{\prime}(s)$ by (5) and perform
$\textstyle\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}(n(s)\hat{V}(s)+n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})$
$\textstyle{}-n(s,a^{i})\hat{Q}(s,a^{i})).$ (13)
3.4 MDP and POMDP Value Gradient
In the previous subsections we’ve assumed that the updated action $a^{\prime}$ is arbitrary. We now present a method to compute an estimate of $\nabla_{a}Q(s,a)$, in order to compute a gradient-based update. Our derivation of $\nabla_{a}Q(s,a)$ is based on the gradient of the transition likelihood, and there are two main advantages to this approach. First, the non-recursive computation can be computed separately for each action branch in the MIS tree, which greatly reduces computational complexity. Second, which is our key observation, is that this derivation extends to POMDPs as well via the propagated belief trick, and has a computable form for the PB-MDP.
Theorem 1.
For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})\textstyle=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\textstyle\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))].$ (14)
Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e. $b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
The gradient is written such that states $s^{\prime}$ could have been sampled from an action $a$ different from $a^{\prime}$. Intuitively, it tends toward actions that both maximize immediate rewards, and maximize the likelihood of future states with high values. Importantly, because $s^{\prime}$ is an integration variable, the future value is not a function of $a^{\prime}$, and has no gradient w.r.t. it.
Theorem 1 is general, and applies to MDPs, state-reward POMDPs, and $\rho$-POMDPs. The following Theorem 2 modifies the gradient for state-reward POMDPs. The gradient for the immediate reward is based on single state samples rather than sampling full beliefs. It may give more accurate gradient estimates, albeit possibly more expensive to compute.
Theorem 2.
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
$\textstyle+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]\textstyle+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\textstyle\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})].$ (15)
For ordered particle beliefs, the gradient log of the propagated transition likelihood from Lemma 1 is given by:
3.5 Action Gradient Monte Carlo Tree Search
Algorithm 1 AGMCTS

### v3

@@ A.5 Derivations of MIS Tree Update Equations
    Proposal transition likelihoods.
    MDP:
    PB-MDP:
@@
    Immediate rewards.
    MDP:
    $\boldsymbol{r}=(r(s,a,s^{\prime,i}))_{i=1}^{M}$ (108)
    PB-MDP:
    $\boldsymbol{r}=(r(\bar{b},a,\bar{b}^{\prime,i}))_{i=1}^{M}.$ (109)
  5. 5.
    Future value estimates.
    MDP:
    $\boldsymbol{v}=(\hat{V}(s^{\prime,i}))_{i=1}^{M}$ (110)
    PB-MDP:
    $\boldsymbol{v}=(\hat{V}(\bar{b}^{\prime,i}))_{i=1}^{M}$ (111)
From here on, the PB-MDP case is the same up to different notations as the MDP case, and we will show only the latter. Using these notations the immediate reward and future value estimates can be compactly written:
$\eta(s,a)=\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\oslash\boldsymbol{q}),$ (112)
$\hat{r}(s,a)=\eta(s,a)^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{r}\oslash\boldsymbol{q}),$ (113)
$\hat{V}_{f}(s,a)=\eta(s,a)^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{v}\oslash\boldsymbol{q}),$ (114)
[…]
We store the following values:
  1. 1.
    $l(\eta)$.
  2. 2.
    Log likelihoods $l(\boldsymbol{p})$ and $l(\boldsymbol{q})$.
  3. 3.
  4. 4.
@@
$\eta^{\prime}(s,a)=\eta(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}\left(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}\right),$ (115)
$\hat{V}_{f}^{\prime}(s,a)=\frac{1}{\eta^{\prime}(s,a)}\left(\eta(s,a)\hat{V}_{f}(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j})\right).$ (116)
Therefore, the log-form update equations are:
$l(\eta^{\prime}(s,a))=LSE\left([l(\eta(s,a)),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})+l(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j})]\right)$ (117)
$\hat{V}_{f}^{\prime}(s,a)=\exp\Big{(}LSE\big{(}[l(\eta(s,a)),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})],[\hat{V}_{f}(s,a),\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}]\big{)}-l(\eta^{\prime}(s,a))\Big{)}.$ (118)
Proof.
We start with the update for $\eta^{\prime}(s,a)$:
$\eta^{\prime}(s,a)-\eta(s,a)$ (119)
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\hat{V}_{f}(s,a)$ (122)
$=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\frac{1}{\eta(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (123)
$=\frac{1}{\eta^{\prime}(s,a)}\left(\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (124)
$=\frac{1}{\eta^{\prime}(s,a)}\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}),$ (125)
@@
Proof.
We append new terms to all vectors: $\boldsymbol{n}_{M+1}$, $\boldsymbol{p}_{M+1}$, $\boldsymbol{v}_{M+1}$, $\boldsymbol{q}_{M+1}$. Update to $\eta(s,a)$:
$\eta^{\prime}(s,a)-\eta(s,a)$ (126)
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\hat{V}_{f}(s,a)$ (129)
$=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\frac{1}{\eta(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (130)
$=\frac{1}{\eta^{\prime}(s,a)}\left(\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (131)
$=\frac{1}{\eta^{\prime}(s,a)}\frac{\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}}\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{v}_{M+1}^{\prime},$ (132)
which again corresponds to (125) with $\boldsymbol{n}_{M+1}=0$. ∎
@@
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega_{a^{\prime}}(s^{\prime,i})$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$\omega_{a^{\prime}}(s^{\prime,i})=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a(s^{\prime,i})))^{-1},$ (133)
and recalculate (100) and (101) with the new $\omega_{a^{\prime}}(s^{\prime,i})$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
Therefore, the log-form equations are:
$l(\eta^{\prime}(s,a))=LSE(l(\boldsymbol{n})\oplus l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}))$ (134)
$\hat{V}_{f}^{\prime}(s,a)=\exp\left(LSE(l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}),\boldsymbol{n}\odot\boldsymbol{v})-l(\eta^{\prime}(s,a))\right)$ (135)
Proof.
$\eta^{\prime}(s,a)=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}}{\boldsymbol{q}_{i}},$ (136)
$\hat{V}_{f}^{\prime}(s,a^{\prime})=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}.$ (137)
@@
Proof.
This is a simple running average update for the mean estimator:
$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (139)
for all $v^{i}$ rollout values from the terminal node $s$. ∎
@@
$\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}\left(n(s)\hat{V}(s)+n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (141)
Proof.
Therefore:
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$ (143)
All terms are equal except for the updated node $j$. Hence, we obtain
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$
Appendix B AGMCTS Algorithm
Algorithm 2 AGMCTS - Full Pseudocode